# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

### v2

@@ B.3.2 Physics-Informed Neural Network (Burgers’ Equation)
The Research Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Research Agent in PDE-constrained optimization. It is tasked with implementing a Python function pinn_loss_strategy(residuals, ic_err, bc_err, extras) that returns a scalar loss tensor. Note that the agent does not solve the optimization problem itself; instead, it defines the scalar objective function that the external optimizer (Adam) minimizes. The prompt explicitly encourages the agent to use adaptive weighting (based on extras[’epoch’]) and to address numerical stiffness (e.g., by re-weighting or normalizing residuals near shocks).
Task-Distributed Multi-LLM Supporters.
@@
The Exploration Supporter (gpt-oss:120b) critiques the generated loss strategy code, identifying issues like scale imbalance between error terms and poor schedule design. For example, it detected that the PDE residual term ($\sim 10^{-2}$) often overshadowed boundary condition errors ($\sim 10^{-5}$), causing ”boundary drift.” In the logs, it explicitly recommended dynamic normalization and component-wise clipping (Huber loss) to fix this: “Replace global clamp with a Huber loss to prevent exploding gradients… Keep a running estimate of the standard deviation of each component and divide by it.” This ensured the optimization respected physical boundaries at $x=\pm 1$.
@@
The Fine-Tuning Strategist (deepseek-r1:32b) dynamically tunes DPO hyperparameters during the periodic fine-tuning updates (every 20 rounds) to ensure the generator learns robust loss structures without collapsing to degenerate solutions. The logs show it actively adjusting $\beta$ (e.g., setting beta=0.7 then 0.5). While it occasionally suggests changes to num_train_epochs or score_threshold (e.g., ”Increasing beta and epochs for better exploration”), the system prioritizes its $\beta$ adjustments to stabilize convergence without inducing mode collapse.
@@
The Policy Inspector (llama4:latest) enforced the improvement threshold. In one instance, it rejected an update where the candidate loss strategy minimized the PDE residual to $10^{-6}$ but caused the Initial Condition error to spike to $10^{-1}$ (a common mode collapse in PINNs known as the ”trivial solution” where $u(t,x)=0$). The inspector flagged this as “Regression: IC error spike” and reverted the reference model, saving the training run from divergence.
B.4 Detailed Agent Dynamics
@@
Task Trial $\Delta\beta$ $\Delta$ Threshold $\Delta$ Epoch Acceptance Rate
Bandit 1 4/25 ($16\%$) 10/25 ($40\%$) 2/25 ($8\%$) 1/25 ($4.0\%$)
2 3/22 ($13.6\%$) 9/22 ($40.9\%$) 2/22 ($9.1\%$) 8/22 ($36.4\%$)
3 7/23 ($30.4\%$) 6/23 ($26.1\%$) 1/23 ($4.3\%$) 1/23 ($4.3\%$)
4 6/22 ($27.3\%$) 6/22 ($27.3\%$) 1/22 ($4.5\%$) 7/22 ($31.8\%$)
5 5/19 ($26.3\%$) 4/19 ($21.1\%$) 2/19 ($10.5\%$) 1/19 ($5.3\%$)
Mean 22.7% 31.1% 7.3% 16.4%
PINN 1 0/10 ($0\%$) 0/10 ($0\%$) 0/10 ($0\%$) 5/10 ($50\%$)
2 3/10 ($30\%$) 4/10 ($40\%$) 1/10 ($10\%$) 7/10 ($70\%$)
3 4/10 ($40\%$) 1/10 ($10\%$) 0/10 ($0\%$) 1/10 ($10\%$)
4 0/10 ($0\%$) 3/10 ($30\%$) 0/10 ($0\%$) 1/10 ($10\%$)
5 1/10 ($10\%$) 4/10 ($40\%$) 0/10 ($0\%$) 5/10 ($50\%$)
Mean 16.0% 24.0% 2.0% 38.0%
Table 5 presents the detailed breakdown of Fine-Tuning Strategist interventions and Policy Inspector decisions across 5 independent trials. The variability in the Bandit task highlights the adaptive nature of the ATLAS framework. For example, the number of active fine-tuning rounds varies across trials (e.g., Trial 1 had 25 rounds, while Trial 5 had 19). This variance occurs because ATLAS employs an efficiency mechanism: if the exploration phase yields no improvement in score, the computationally expensive fine-tuning loop is skipped for that round. In contrast, the PINN task consistently achieved score improvements, triggering the fine-tuning loop in all 10 scheduled phases across all trials.
B.5 Hyperparameters, Hardware, and Timing
@@
Number of Seeds 55 (Independent Runs)
@@
Initial Beta ($\beta$) $0.6$
@@
Hardware. For benchmarking each task, we were using 8 Nvidia H100 GPUs with 80 GB of RAM. From this cluster, we specifically allocated our developing research agent (Llama-3.2-1B-Instruct), which used 1 Nvidia H100 GPU with 40 GB of RAM and 14 CPU cores.
@@
B.6 KL Proxy and Limitations
@@ B.8 Implementation Details
Appendix C Prompt Templates for Code Generation
This section provides templates for the given prompts for LLM agents. All tasks use the same Fine-Tuning Strategist and Policy Inspector prompt, whereas Exploration Supporter and Self-Evolving Research Agent use suitable prompts for each task.
C.1 Problem Description: Non-stationary Contextual Bandit
@@
- policy.lambda_reg (ridge regularizer, float >> 0)
- policy.window_size (sliding window size W, int >>= 1)
C.2 Problem Description: PINN Burgers
We are evolving loss functions for a physics-informed neural network (PINN) that solves the 1D viscous Burgers equation: u_t + u * u_x - nu * u_xx = 0, (t, x) in [0, 1] x [-1, 1] with:
@@ C.2 Problem Description: PINN Burgers
The strategy must combine these into a scalar loss tensor. During evaluation, we compute a scalar total_err = pinn_loss_strategy(...) on held-out samples, then convert it to a scalar score:
err_mag = abs(total_err)
score = -err_mag
So:
- Smaller |total_err| =>> larger score (closer to 0, still negative).
- Larger |total_err| =>> more negative score.
The goal is to design loss strategies that make the PINN satisfy the PDE, initial condition, and boundary conditions as accurately and robustly as possible.
C.3 Self-Evolving Research Agent: Non-stationary Contextual Bandit
You are an research agent in online learning and contextual bandits.
@@ C.3 Self-Evolving Research Agent: Non-stationary Contextual Bandit
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) ->> int
@@ 8. You tune hyperparameters using feedback from teacher:
policy.window_size (int >>= 1)
policy.lambda_reg (float >> 0)
policy.ucb_alpha (float >>= 0)
@@
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) ->> int
@@
C.4 Self-Evolving Research Agent: PINN Burger’s equation
@@
- Example: return residuals.mean() + ic_err.mean() ->> CORRECT
- Example: return lambda: residuals.mean() ->> WRONG (returns a function)
@@
C.5 Exploration Supporter: Non-stationary Contextual Bandit

### v3

@@ C.3.2 Physics-Informed Neural Network (Burgers’ Equation)
The Evolving Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Evolving Agent in PDE-constrained optimization. It is tasked with implementing a Python function pinn_loss_strategy(residuals, ic_err, bc_err, extras) that returns a scalar loss tensor. Note that the agent does not solve the optimization problem itself; instead, it defines the scalar objective function that the external optimizer (Adam) minimizes. The prompt explicitly encourages the agent to use adaptive weighting (based on extras[’epoch’]) and to address numerical stiffness (e.g., by re-weighting or normalizing residuals near shocks).
Role-Specialized Multi-LLM Training Supporters.
@@
The Exploration Supporter (gpt-oss:120b) critiques the generated loss strategy code, identifying issues like scale imbalance between error terms and poor schedule design. For example, it detected that the PDE residual term ($\sim 10^{-2}$) often overshadowed boundary condition errors ($\sim 10^{-5}$), causing "boundary drift." In the logs, it explicitly recommended dynamic normalization and component-wise clipping (Huber loss) to fix this: “Replace global clamp with a Huber loss to prevent exploding gradients… Keep a running estimate of the standard deviation of each component and divide by it.” This ensured the optimization respected physical boundaries at $x=\pm 1$.
@@
The Fine-Tuning Strategist (Qwen 3-32b) dynamically tunes DPO hyperparameters during the periodic fine-tuning updates (every 20 rounds) to ensure the generator learns robust loss structures without collapsing to degenerate solutions. The logs show it actively adjusting $\beta_{DPO}$ (e.g., setting beta=0.7 then 0.5). While it occasionally suggests changes to num_train_epochs or score_threshold (e.g., "Increasing beta and epochs for better exploration"), the system prioritizes its $\beta_{DPO}$ adjustments to stabilize convergence without inducing mode collapse.
@@
The Policy Inspector (llama4:latest) enforced the improvement threshold. In one instance, it rejected an update where the candidate loss strategy minimized the PDE residual to $10^{-6}$ but caused the Initial Condition error to spike to $10^{-1}$ (a common mode collapse in PINNs known as the "trivial solution" where $u(t,x)=0$). The inspector flagged this as “Regression: IC error spike” and reverted the reference model, saving the training run from divergence.
C.3.3 Traveling Salesperson Problem (TSP)
Evolving Agent.
The Evolving Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Evolving Agent for combinatorial optimization. For each TSP instance, the policy receives a node set and a distance matrix, and it generates a structured JSON output containing a candidate tour. The generated tour is then passed to a deterministic verifier, which checks whether the tour visits every node exactly once and returns to the starting node. The verifier also computes the tour length and the optimality gap relative to the exact-solver reference. Thus, the Evolving Agent is not rewarded for self-reported reasoning quality; it is scored only through executable validity and objective value.
Role-Specialized Multi-LLM Training Supporters.
The Exploration Supporter (gpt-oss:120b) analyzes generated TSP heuristics and proposes structural improvements to the candidate-generation strategy. For example, when the Evolving Agent repeatedly produced nearest-neighbor-style tours, the supporter encouraged additional local-improvement operations such as 2-opt swaps, edge-crossing removal, and route repair for duplicated or omitted nodes. These patches expand the candidate distribution beyond purely greedy construction and help prevent the policy from converging to a narrow heuristic family.
The Fine-Tuning Strategist (Qwen 3-32b) monitors the distribution of verifier scores and adjusts the DPO fine-tuning hyperparameters, including $\beta_{\mathrm{DPO}}$, score-margin thresholds, and training epochs. In TSP, small score differences can arise from minor edge substitutions, while invalid tours should be strongly separated from valid but suboptimal tours. The strategist therefore controls the score threshold used for preference-pair construction so that the DPO update emphasizes meaningful improvements in tour quality rather than noisy or near-equivalent route variations.
The Policy Inspector (llama4:latest) acts as the reference-promotion gate for candidate TSP policies. It accepts a proposed reference only when the candidate improves verifier-based tour quality while satisfying the trajectory-conditioned proxy-KL constraint. This is important because some candidate policies can improve a small set of TSP instances by making aggressive heuristic changes, but degrade validity or generalization on other instances. The inspector therefore prevents high-variance route-generation strategies from immediately replacing the stable reference policy.
C.3.4 Bin Packing Problem
Evolving Agent.
The Evolving Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Evolving Agent for discrete resource-allocation problems. For each Bin Packing instance, the policy receives item weights and a bin capacity, and it generates a structured JSON output representing an assignment of items to bins. A deterministic verifier checks that every item is assigned exactly once, no item is duplicated, and no bin exceeds the capacity constraint. The verifier then computes the number of bins used, extra bins relative to the exact integer-programming reference, and packing density. Candidate preferences are induced from this verifier score rather than from model-generated judgments.
Role-Specialized Multi-LLM Training Supporters.
The Exploration Supporter (gpt-oss:120b) critiques generated packing heuristics and proposes structural modifications to improve feasibility and bin utilization. For example, when the Evolving Agent relied too heavily on a simple first-fit or first-fit-decreasing rule, the supporter suggested refinements such as residual-capacity tracking, best-fit insertion, repair of overloaded bins, and reassignment of small items to improve density. These modifications help the policy escape low-density packings and reduce unnecessary bin usage.
The Fine-Tuning Strategist (Qwen 3-32b) adjusts DPO hyperparameters based on verifier-score distributions across candidate packings. In Bin Packing, many candidate outputs may be feasible but differ only slightly in density, while infeasible outputs with capacity violations must be strongly penalized. The strategist therefore modulates $\beta_{\mathrm{DPO}}$ and score thresholds to emphasize preference pairs that separate feasible, compact packings from infeasible or inefficient assignments. This prevents the Evolving Agent from overfitting to weak preferences generated by nearly equivalent packings.
The Policy Inspector (llama4:latest) enforces conservative reference promotion for Bin Packing. Because small changes in assignment logic can produce large feasibility shifts, the inspector promotes a candidate reference only if it improves verifier-based packing quality while remaining within the proxy-KL budget. In practice, this makes the inspector a safeguard against unstable heuristic changes, such as policies that reduce the number of bins on some instances but increase capacity violations or unassigned-item errors on others.
C.4 Detailed Agent Dynamics
Table 4 summarizes the intervention frequencies of the Fine-Tuning Strategist and the acceptance behavior of the Policy Inspector across 10 independent seeds. The intervention rates are computed over active fine-tuning rounds, i.e., rounds in which the exploration stage produced candidates that triggered a fine-tuning update. Overall, the Strategist intervenes frequently across all domains, but the type of intervention varies by task. Epoch adjustment is triggered in nearly all active fine-tuning rounds across the four domains, suggesting that the training horizon required continuous adaptation as candidate quality and score distributions changed. In contrast, the $\beta_{\mathrm{DPO}}$ and score-threshold interventions are more task-dependent: Bin Packing requires the most frequent $\beta_{\mathrm{DPO}}$ adjustment, while PINN and TSP require especially frequent threshold adjustment.
The Policy Inspector shows a different pattern. Bandit has the highest reference-promotion acceptance rate ($67.5\%$), indicating that candidate references often improved evaluator score while remaining within the proxy-KL budget. PINN has a moderate acceptance rate ($37.0\%$), consistent with smoother but still selective reference progression. By contrast, TSP and Bin Packing have much lower acceptance rates ($4.0\%$ and $7.7\%$, respectively), suggesting that combinatorial heuristic updates are more volatile: proposed candidate references may improve isolated instances but often fail the score-improvement and proxy-KL gate required for reference promotion. This behavior supports the intended role of the Policy Inspector as a conservative safeguard rather than an automatic reference-refresh mechanism.
@@ C.4 Detailed Agent Dynamics
Task seeds $\Delta\beta_{\mathrm{DPO}}$ $\Delta$ Threshold $\Delta$ Epoch Acceptance Rate
Bandit 10 67.5% 49.1% 99.1% 67.5%
PINN 10 49.5% 88.7% 99.0% 37.0%
TSP 10 40.7% 91.9% 97.2% 4.0%
Bin Packing 10 92.0% 35.9% 95.0% 7.7%
Table 4 also highlights that ATLAS does not apply the same control policy uniformly across domains. In the Bandit task, both $\beta_{\mathrm{DPO}}$ adjustment and reference promotion are frequent, reflecting the need to track a drifting reward model while still allowing stable reference progression. In PINN, the high threshold-adjustment rate suggests that preference-pair selection must be actively controlled as the optimization landscape becomes stiff. In TSP and Bin Packing, the low acceptance rates indicate that the inspector rejects most candidate references, which is expected when discrete heuristic changes cause high-variance performance shifts across instances.
The number of active fine-tuning rounds may vary across seeds because ATLAS skips expensive fine-tuning when the exploration phase fails to produce candidates with sufficient evaluator-score improvement. This mechanism reduces unnecessary updates and prevents the preference dataset from being dominated by low-quality or uninformative candidate pairs. Consequently, the reported intervention rates should be interpreted as conditional rates over active fine-tuning rounds rather than over all exploration rounds.
C.5 Hyperparameters, Hardware, and Timing
@@
Number of Seeds $10$ (Independent Runs)
@@
Initial Beta ($\beta_{DPO}$) $0.6$
@@
TSP Task Instance Generation Uniform nodes in $0,1]^{2}$
Problem Sizes ($n$) $10,20,50,100$ nodes
Distance Metric Euclidean distance
Output Format JSON tour
Maximum Refinement Rounds 55
Exact Evaluation Baseline Exact TSP solver
Primary Metric Relative improvement
Bin Packing Task Instance Generation Random item weights
Problem Sizes ($n$) $10,50,100$ items
Bin Capacity ($C$) Fixed per instance
Output Format JSON bin assignment
Maximum Refinement Rounds 55
Exact Evaluation Baseline Integer-programming solver
Primary Metric Relative improvement
@@
Hardware. For benchmarking each task, we were using 8 Nvidia l40s GPUs with 20 GB of RAM. From this cluster, we specifically allocated our Evolving Agent (Llama-3.2-1B-Instruct), which used 1 Nvidia l40s GPU with 20 GB of RAM and 12 CPU cores.
@@
External assets, models, and solvers.
All experiments use publicly described or API-accessed LLM backbones and standard optimization or evaluation software. We do not redistribute model weights, third-party solver code, or proprietary API assets. The LLM backbones used in our experiments are Llama-3.2-1B-Instruct for the Evolving Agent, gpt-oss-120B for the Exploration Supporter, Qwen 3-32b for the Fine-Tuning Strategist, and Llama4-latest for the Policy Inspector. For combinatorial evaluation, exact or integer-programming solvers are used only to compute reference objective values and feasibility checks; they do not provide candidate solutions to ATLAS. These external tools are used under their respective access terms and are listed here solely for reproducibility.
C.6 KL Proxy and Limitations
@@ C.8 Implementation Details
Appendix D Limitations, Broader Impact, and Ethics
Limitations.
Our theoretical analysis is intentionally stylized and isolates adaptive reference progression in a non-stationary preference contextual bandit. It does not provide an end-to-end guarantee for full sequence-level LLM fine-tuning, and it does not directly model the finite checkpoint candidate set, the empirical token-level KL proxy, or the discrete accept/reject behavior of the inspector gate. Empirically, our experiments focus on four executable optimization domains: non-stationary contextual bandits, PINN loss reweighting, TSP, and Bin Packing. Although these tasks cover decision-making, scientific machine learning, and combinatorial optimization, they do not exhaustively represent all open-ended scientific-discovery settings.
Broader impact.
ATLAS is intended to improve evaluator-driven self-improvement for executable optimization tasks, including scientific computing and algorithmic search. Potential positive impacts include reducing manual effort in designing optimization heuristics, stabilizing long-horizon preference-based training, and improving scientific-computing workflows where evaluator feedback is available. Potential negative impacts include over-reliance on automatically generated code, propagation of evaluator misspecification, generation of incorrect or inefficient programs, and misuse of autonomous code-improvement pipelines in settings where generated code could affect safety-critical systems. We mitigate these risks in our experiments by using task-specific executable evaluators, fixed external validation metrics for PINN evaluation, and a conservative KL-based reference-promotion gate. The proposed system should not be deployed in safety-critical settings without human review, stronger sandboxing, and domain-specific validation.
Ethics and data.
The experiments do not involve human subjects, private personal data, crowdsourcing, or scraped user datasets. All preference labels are induced by executable task evaluators rather than human annotators. The work uses existing LLMs and software frameworks; their licenses and access terms should be respected when reproducing or extending the system.
Appendix E Prompt Templates for Code Generation
This section provides templates for the given prompts for LLM agents. All tasks use the same Fine-Tuning Strategist and Policy Inspector prompt, whereas Exploration Supporter and Evolving Agent use suitable prompts for each task.
E.1 Problem Description: Non-stationary Contextual Bandit
@@
- policy.lambda_reg (ridge regularizer, float > 0)
- policy.window_size (sliding window size W, int >= 1)
E.2 Problem Description: PINN Burgers
We are evolving training loss functions for a physics-informed neural network (PINN) that solves the 1D viscous Burgers equation: u_t + u * u_x - nu * u_xx = 0, (t, x) in [0, 1] x [-1, 1] with:
@@ E.2 Problem Description: PINN Burgers
The candidate function must combine these quantities into a scalar training loss tensor used only for inner PINN optimization.
Candidate-generated losses are not used directly as the evaluation score. After training a PINN with the candidate loss, an external fixed evaluator computes validation error using evaluator-defined terms:
- normalized PDE residual error,
- initial-condition error,
- boundary-condition error,
- solution L2 error against a high-resolution reference solution.
The scalar score is the negative of this fixed evaluator error:
score = -fixed_evaluator_error
Thus:
- Lower fixed evaluator error gives a larger score.
- Higher fixed evaluator error gives a smaller score.
The goal is to design training loss strategies that lead to low fixed-evaluator error after PINN optimization. The generated loss may shape training, but all preference labels and final metrics are computed using the fixed external evaluator.
E.3 Problem Description: Traveling Salesperson Problem
Traveling Salesperson Problem (TSP): You are given n nodes and a distance matrix D.
D[i][j] is the distance from node i to node j.
Goal:
Find a valid tour that:
- starts at node 0,
- visits every node exactly once,
- returns to node 0,
- minimizes the total tour length.
A candidate tour is evaluated by an external verifier. The verifier checks:
- whether every node is visited exactly once,
- whether the tour starts and ends at node 0,
- whether the output is valid JSON,
- the total tour length.
The score is based on the tour length and feasibility. Invalid tours are penalized.
The model must output a machine-readable JSON object so the verifier can parse it.
E.4 Problem Description: Bin Packing
Bin Packing Problem (BPP): You are given n items with positive weights and a fixed bin capacity C.
Goal:
Assign every item to a bin while:
- using as few bins as possible,
- never exceeding the capacity C in any bin,
- assigning every item exactly once,
- avoiding duplicated or missing items.
A candidate packing is evaluated by an external verifier. The verifier checks:
- whether every item is assigned exactly once,
- whether any bin exceeds capacity,
- whether the output is valid JSON,
- the number of bins used,
- the packing density.
The score is based on feasibility and the number of bins used. Invalid packings are penalized.
The model must output a machine-readable JSON object so the verifier can parse it.
E.5 Evolving Agent: Non-stationary Contextual Bandit
You are an Evolving Agent in online learning and contextual bandits.
@@ E.5 Evolving Agent: Non-stationary Contextual Bandit
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) -> int
@@ 8. You tune hyperparameters using feedback from teacher:
policy.window_size (int >= 1)
policy.lambda_reg (float > 0)
policy.ucb_alpha (float >= 0)
@@
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) -> int
@@
E.6 Evolving Agent: PINN Burger’s equation
@@
- Example: return residuals.mean() + ic_err.mean() -> CORRECT
- Example: return lambda: residuals.mean() -> WRONG (returns a function)
@@
E.7 Evolving Agent: Traveling Salesperson Problem
You are an Evolving Agent for combinatorial optimization.
TASK:
Generate a valid Traveling Salesperson Problem (TSP) tour.
You are given:
- n: number of nodes
- distance_matrix: a 2D list or array where distance_matrix[i][j] is the distance from node i to node j
Output:
Return ONLY a valid JSON object with the following structure:
"tour": [0, ..., 0]
Rules:
1. The tour MUST start at node 0 and end at node 0.
2. Every node from 0 to n-1 MUST appear exactly once before returning to node 0.
3. Do NOT duplicate nodes except for the final return to node 0.
4. Do NOT omit any node.
5. Do NOT include nodes outside the range [0, n-1].
6. The output MUST be valid JSON.
7. Do NOT include markdown, comments, or extra text outside the JSON object.
8. The evaluator will compute the total tour length. Do NOT self-report the score.
Goal:
Produce a feasible tour with small total distance.
You may use constructive heuristics such as nearest neighbor, insertion, or local repair,
but the final answer must be only the JSON object.
E.8 Evolving Agent: Bin Packing
You are an Evolving Agent for discrete resource allocation.
TASK:
Generate a feasible bin packing assignment.
You are given:
- capacity: maximum capacity of each bin
- item_weights: a list of positive item weights
Output:
Return ONLY a valid JSON object with the following structure:
"bins": $[$...], [...], ...]
Each inner list contains item indices assigned to that bin.
Rules:
1. Every item index from 0 to len(item_weights)-1 MUST appear exactly once.
2. Do NOT duplicate any item.
3. Do NOT omit any item.
4. The total weight in each bin MUST be less than or equal to capacity.
5. Do NOT include item indices outside the valid range.
6. The output MUST be valid JSON.
7. Do NOT include markdown, comments, or extra text outside the JSON object.
8. The evaluator will compute feasibility, number of bins, and packing density. Do NOT self-report the score.
Goal:
Use as few bins as possible while satisfying all capacity constraints.
You may use constructive heuristics such as first-fit decreasing, best-fit decreasing,
or local repair, but the final answer must be only the JSON object.
E.9 Exploration Supporter: Non-stationary Contextual Bandit